OpenStudy (cwrw238):

Just for fun . A quiz question: If H is 361 and I is 529 what is A?

OpenStudy (anonymous):

playin around AP or GP stuff?

OpenStudy (ghazi):

248

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

no - need a hint prime numbers

OpenStudy (anonymous):

4

OpenStudy (ghazi):

19*19=361, 23*23=529 so A must be 7*7=49 or 5*5=25

OpenStudy (cwrw238):

right idea ghazi but its 2^2 = 4

OpenStudy (ghazi):

oh missed it.. :)